OpenStudy (konradzuse):

Find the indefinite integral sec^2 and tan

OpenStudy (konradzuse):

\[\int\limits \sqrt{\tan(3x)} * (\sec(3x))^2 dx\]

OpenStudy (konradzuse):

I figured I'd do u = sec(3x)^2 du = tan(3x) But that sqrt is in the way... I cannot have a sqrt(du) right?

OpenStudy (anonymous):

try \(u=\tan(3x)\) to get what you want almost immediately

OpenStudy (konradzuse):

oh wait I did it backwards LOL.

OpenStudy (anonymous):

you have it backwards derivative of tangent is secant squared

OpenStudy (konradzuse):

There is no derivative of sec^2 :P too tired :P

OpenStudy (konradzuse):

so u = tan(3x) du = sec(3x)^2 Does the chain rule apply here?

OpenStudy (konradzuse):

sec(3x)^2 dx

OpenStudy (konradzuse):

so then I have \[\sqrt{u}du --- u^{1/2} du\]

OpenStudy (konradzuse):

\[\frac{2(u)^{3/2}}{3} + c\]

OpenStudy (anonymous):

hold on a sec

OpenStudy (konradzuse):

\[\frac{2(tan(3x))^{3/2}}{3} + c\]

OpenStudy (anonymous):

\[u=\tan(3x),du=3\sec^2(3x)dx\] by the chain rule so \(\frac{1}{3}du=\sec^2(3x)dx\)

OpenStudy (konradzuse):

Ok so we do use the chain rule, ty.

OpenStudy (konradzuse):

so the end should be 2/6 then?

OpenStudy (anonymous):

so you have \[\frac{1}{3}\int\sqrt{u}du\] etc

OpenStudy (anonymous):

anti derivative of \(\sqrt{u}\) is \(\frac{2}{3}u^{\frac{3}{2}}\) so you get \[\frac{2}{9}u^{\frac{3}{2}}\]

OpenStudy (anonymous):

yeah it is \(3\times 3=9\) and not 6 right? take a break it is saturday night have a beer or whatever

OpenStudy (konradzuse):

I thought it was 1/3 * 1/2 ooopsies :P

OpenStudy (konradzuse):

Doing things too fast :P

OpenStudy (konradzuse):

\[\frac{2(\tan(3x))^{3/2}}{9} + c\]